OpenStudy (anonymous):

Medal to whoever provides correct answer or helps me A bacteria culture begins with 4 bacteria which double in size every hour. How many bacteria exist in the culture after 8 hours? 128 256 1,024 8,192

OpenStudy (anonymous):

1,024 :)

OpenStudy (zolock):

1) y=4(2)^n Let n = 8 y = 4(2)^8 y = 4(256) y = 1024 The number of bacteria will be 1024 after 8 hours.

OpenStudy (anonymous):

4x2^8= 1024 <---

OpenStudy (anonymous):

\[4 \times 2^{8}\] = 1024.